OpenStudy (tiffclifford13):

The intercepts of the equation x^2 + 4y^2 = 16 are ___ Please help!

OpenStudy (wolf1728):

That is a quadratic equation. Do you know how to solve it?

OpenStudy (tiffclifford13):

I'm not sure I am so bad at Trig would you use the quadratic formula?

OpenStudy (wolf1728):

oops I'm sorry that is a quadratic equation in TWO unknowns geez - that makes it a lot tougher.

OpenStudy (tiffclifford13):

do you at least have a guess of what it would be im so lost

OpenStudy (wolf1728):

x^2 + 4y^2 = 16 It could be rewritten -4y^2 = x^2 -16 -4y^2 = (X+1)*(X-1) that's still not a solution

OpenStudy (tiffclifford13):

okay thanks for the help though i appreciate it

OpenStudy (wolf1728):

sorry :-(

OpenStudy (ajspeller):

Subtract 16 from both sides and then factor. Please take a look at my various tutorials on the subject. They are very easy to follow and short. They also explain each detail and lastly they show how to use a graphing calculator to help you as well. I hope these help. https://www.youtube.com/playlist?list=PLsX0tNIJwRTxqIIcfpsTII9_xhqVIbQlp